# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ III-E Threat models
The previous taxonomy with five dimensions provides us with a powerful tool for grouping and classifying FA algorithms and systems. In Table III, we list some existing FA solutions, and show their positions according to our taxonomy. [^1] Throughout the survey, the client scale, or maximal supported number of clients, is counted as the maximal number of clients in these papers’ experiment settings.
@@ IV Key challenges in Federated Analytics
IV-A Accuracy, privacy, and performance trade-off
@@
FA aims to extract and analyze insights from decentralized data sources in a privacy-preserving way. There are trade-offs as regards accuracy, privacy, and performance, which are three of the most important aspects of FA. Researchers in [49] pointed out these trade-offs among differential privacy, secure multi-party computation, and approximate query processing. Firstly, when discussing trade-offs between accuracy and privacy, it is imperative to bring up the differential privacy (DP) which has been a widely acknowledged privacy-preserving criterion in data analytics. The details of DP are described in Section V-A herein. To satisfy DP, mechanisms usually introduce random noise into a result. An increase in noise addition corresponds to heightened privacy protection since it more and more obfuscates the contribution of each individual. However, the guarantee of privacy comes at the cost of a reduction in accuracy, and the result deviates from the ground truth due to the injection of noise. Secondly, the trade-off including performance is usually related to another sort of privatization technique related to cryptography. Although it guarantees both privacy and accuracy by encrypting the data of each individual before sharing it and obtaining the real result by corresponding decryption, there is an enormous amount of computation and communication overhead when performing such operations in FA. Lastly, there are approximating methods applied in FA to increase the performance of the system which sacrifices the accuracy of results. For example, private sampling algorithms utilize less data in the computation and communication process while introducing variance from sampling into the final result. In conclusion, balancing these trade-offs is a key challenge in FA.
IV-B Intensive computation
@@
In FA, cryptography is commonly used to protect the privacy of each client’s data while it is processed and analyzed by multiple parties including the aggregator and other clients. However, cryptographic methods are considerably resource-intensive, both in terms of computation and communication. Operations in cryptography like encryption, decryption, and hashing are computationally intensive. For some advanced cryptographic techniques used in FA like homomorphic encryption and secure multi-party computation which allow various computations on the encrypted data, the required computation resource is even higher [65]. Regarding communication cost, encrypted data tends to be larger than their plaintext counterparts. Furthermore, most cryptographic protocols require multiple rounds of communication between parties [10], exchanging messages like intermediate results. This attribute of cryptography extremely increases the bandwidth required for tasks in FA. In addition, managing cryptographic keys is also a complex task in federated settings. They are required to be private and secure in their generation, distribution, rotation, and destruction [66]. This process will be more complex and computationally intensive if the number of clients increases [29]. These attributes together contribute to the large resource demands for utilizing cryptography in FA, which is a challenge to preserve privacy efficiently in terms of computation costs.
IV-C Wireless communication
@@
Wireless communication becomes a major challenge in FA primarily due to its distributed and dynamic nature. The slowest participant in this networked system can significantly hinder overall performance, known as the “straggler effect”. This is often exacerbated by varying data transmission rates and unreliable communication channels, leading to delays and inconsistencies in data aggregation. The paper [67] addresses it in FL by clustering user equipment based on upload times, thereby reducing time divergence among participants. In addition, the over-the-air (OTA) [68] technique brings efficiency as well as new challenges to federated analytics and federated learning, such as signal interference and the requirement for precise synchronization. Furthermore, the mobility of devices in wireless networks, such as in vehicular networks and aerial networks, introduces another layer of complexity, requiring adaptive strategies for data transmission and model training that can cope with changing network typologies and varying channel conditions. The integration of technologies like Reconfigurable Intelligent Surfaces (RIS) shows promise in enhancing communication efficiency and reliability in these scenarios [69], while they also require sophisticated optimization of resources and careful handling of estimated Channel State Information (CSI).
IV-D Limited privacy resource
@@
Privacy resource in the context of FA usually refers to the “privacy budget” in the DP mechanisms, which intuitively means the amount of allowable privacy loss for each individual within an analysis of the whole data set. Although the information of clients is protected by various privacy-preserving techniques, there is still a risk that repeated analyses such as multiple or iterative queries can expose it. Typically, a portion of the privacy budget is consumed and a specific amount of noise will be added to the results for a query. Managing the privacy budget is important since multiple analyses may be performed on the same group of clients. Every analysis incrementally raises the risk of privacy loss so this cumulative effect ought to be monitored and controlled. How to effectively schedule the privacy budget for the multi-query analysis is a challenge in FA concerning the accuracy of results for each query and total privacy loss for each client.
IV-E Robustness
The robustness challenge in FA encompasses a range of considerations that are similar in FL, including malicious and non-malicious data corruption. Malicious data corruption involving deliberate attacks by malicious parties is a considerable threat to the integrity and security of FA systems. These attacks such as data poisoning and privacy breaches call for corresponding robust security measures, such as encryption and authentication to discover and defend. On the other hand, non-malicious data corruption caused by unintentional errors and client dropout due to either connection issues or device failure introduces bias to influence the accuracy of analytics or even causes failure of the system. In this context, there’s a requirement for fault-tolerant mechanisms to cope with client dropouts for robustness, ensuring that the system works effectively even with such disruptions. Addressing these challenges is crucial to maintaining the robustness of FA systems in terms of reliability, security, and accuracy.
IV-F Data heterogenity
@@
Data heterogeneity, i.e., the phenomenon that data possessed by different clients follows heterogeneous distributions, is common in federated systems. As the tenet of privacy preservation restricts raw data from being transmitted, data reallocation or adjustment is usually unavailable for federated systems, making data heterogeneity unavoidable. The challenge of data heterogeneity has become one of the most popular research focuses in the field of FL [70], and also widely exist in FA systems. Since FA systems tackle various data analytics tasks, data heterogeneity introduces diverse influences for different heterogeneous tasks. In the federated MAB model [59, 60, 62], data heterogeneity is formulated as the biased reward of the same arm from different clients, and FA schemes are proposed to guarantee the performance of the model globally or personally. In [17], data heterogeneity is studied in a federated clustering problem, where the data points held by one client are biased to one or several clusters. The authors of [17] found that the data heterogeneity is beneficial for the FA scheme, where a higher analytics performance can be achieved in the heterogeneous data environment. In [8], an FA scheme is proposed to measure the severity of data heterogeneity directly. The proposed scheme can figure out clients with low/high data heterogeneity, which provides significant information to other federated tasks in handling the data heterogeneity challenges.
IV-G Fairness
The challenge of fairness has been investigated in many federated systems, mostly for FL. As the term “fairness” is a comprehensive term. Researchers investigate multiple interpretations of fairness, based on different types of benefits focused by the players. The existing works mainly consider three dimensions of fairness. 1) Collaboration fairness, where some monetary rewards or their equivalents are assumed to be awarded to the participating clients after the federated model is trained, and the participating clients would like to fairly allocate such benefits based on their contribution to the federated system [71, 72, 73, 74]; 2) Model performance fairness, where the finally derived federated model is adjusted to obtain a fair and balanced performance among the clients with heterogeneous data [75, 76, 77]; 3) Participation fairness, which focuses on balancing the chance of clients to become participants of the federated system [78].
Although existing fair federated system research mainly focuses on FL, the rationale and fairness concepts are still applicable to FA algorithms. In addition, the novel characteristics of FA, including the vast investigated data analytics problems, may raise new fairness concepts and challenges, which require tailored solution concepts.
@@ V Key Techniques in Federated Analytics
In this section, we investigate the key enabling techniques that are frequently utilized by FA. These key techniques are categorized by their utilization in FA: In Section V-A, the privatization techniques, i.e., those enhance the privacy preservation of FA, are discussed; in Section V-B, the analytics techniques, i.e., those help derive the analytics results, are presented; in Section V-C, the deployment and optimization techniques, which improves the system availability, are presented
@@ V-A Privatization techniques
The tenet of FA prevents the exposure of raw data to any entity other than the data owner (client), which naturally provides privacy preservation to some extent. However, such a “hiding raw data” idea fails to provide any privacy guarantee. Therefore, existing FA research prefers to enhance the privacy preservation of their schemes by providing some formal and rigid privacy guarantee, by leveraging various privatization techniques. In this part, we analyze the privacy guarantees and the corresponding techniques to realize these techniques.
@@
Local differential privacy Differential privacy (DP), introduced by [79], is a privatization technique enforced on any data publication mechanism. It has become the most popular criterion for privacy preservation in data analytics, as well as FA. It considers the threat models where the data publication mechanism publishes some outputs based on the raw data, and an adversary tries to infer sensitive information based on the output, such as membership inference attacks [80] and data reconstruction attacks [81]. DP handles such threat models by introducing randomness to the output of the data publication mechanism, and restricting the probabilistic distribution of outputs.
@@
Local differential privacy (LDP) is the most widely used DP variation in FA, as it provides the strong local privacy preservation required by the majority of FA applications. LDP-based FA schemes consider the threat model that the clients upload their insights to an untrusted server, where the untrusted server directly leverages the received insights to perform inferences. LDP is enforced by applying randomness in client uploads, so that the probability of any output would not change intensely for clients with arbitrary different local data. The definition of LDP in an FA setting is given as follows.
@@
The satisfaction of LDP requires perturbation (randomness) on the client uploads. For the binary data, LDP is typically realized by the randomized response, i.e., flipping each bit with some probability. The randomized response-based LDP mechanisms usually work in FA applications where the local computation results are naturally in the binary form [21, 16, 38, 40] or the client data are encoded into a binary form with sketching techniques [15, 42, 82]. For the continuous scaler data, LDP is typically realized by introducing continuous scaler noise on the data. Gaussian mechanism and Laplace mechanism are the most widely used schemes to realize LDP for continuous scaler data. Gaussian mechanism is applied by adding a random noise following the Gaussian distribution into the uploaded scaler, which can guarantee the corresponding $(\epsilon,\delta)$-LDP [83, 84]. Laplace mechanism, introducing a random noise following the Laplace distribution, is able to satisfy the stronger $\epsilon$-LDP, which have been utilized in [23, 85, 55].
Central differential privacy Central differential privacy (CDP), also known as global differential privacy, is the original form of DP. It considers the scenario that a data analytics mechanism processes a database containing many data samples, and the published analytics results can protect the raw data of any single data sample from being inferred by an adversary. The definition of CDP is as follows.

### v3

@@ III-D5 Privatization methodology
The previous taxonomy with five dimensions provides us with a powerful tool for grouping and classifying FA algorithms and systems. In Table IV, we list all FA solutions covered in this survey and show their positions according to our taxonomy. [^2] If an FA system does not explicitly clarify its cross-silo FA or cross-device FA, we classify it based on its scale of clients.
@@ IV Key challenges in Federated Analytics
Some of these challenges are unique to FA, such as diverse tasks in terms of lack of a unified framework, defining proper insight forms and aggregation methods, and handling scalability at an unprecedented level. While others are shared with FL, primarily due to the common principles of data decentralization and privacy preservation. However, these common challenges take on distinct forms in FA. Unlike FL, where aggregation primarily deals with model updates, FA requires task-specific aggregation strategies across diverse analytics tasks, making it nontrivial to standardize the framework. Moreover, FA systems often scale to millions or even billions of clients, where every device can contribute to the final analytics results—introducing new challenges in selecting representative clients to ensure statistical reliability. This contrasts with FL, where training a general model can rely on virtual representative datasets obtained via client sampling. Additionally, while avoiding raw data transmission is already a strong privacy guarantee in FL, FA imposes stricter privacy requirements, as descriptive analytics often involve direct statistical summaries. This calls for a more careful trade-off between privacy and utility, further complicated by the scalability of FA systems, making privacy-preserving data aggregation substantially more challenging than in FL.
IV-A Diversity of analytics task
A significant challenge unique to FA lies in the diversity of analytics tasks, which necessitates handling a wide array of data structures, insight extraction methods, and result formats—elements not typically encountered in FL. In FA, tasks vary not only in their objectives but also in the types of input data they handle. Also, this variability requires that FA systems be highly versatile in processing and integrating such diverse data forms to maintain accuracy and consistency in insight extraction. Besides, each analytics task requires specialized aggregation methods designed for the specific type of insights. Furthermore, the output from these tasks also presents varied data structures, from simple numeric aggregates to intricate data visualizations. These aspects significantly complicate the development of universal FA solutions, necessitating tailored approaches that can adapt to the specific requirements of each analytics task while upholding stringent standards of data privacy and system efficiency. Since FA schemes tackling different data analytics tasks behave quite differently regarding key characteristics. The host data analytics task becomes a good indicator in the taxonomy of FA, as we state in Section III-D.
IV-B Accuracy, privacy, and performance trade-off
@@
FA aims to extract and analyze insights from decentralized data sources in a privacy-preserving way. There are trade-offs as regards accuracy, privacy, and performance, which are three of the most important aspects of FA. Firstly, FA schemes utilizing the widely acknowledged privacy-preserving criterion – DP mechanisms, naturally introduce the trade-off between accuracy and privacy. The details of DP are described in Section V-A herein. DP mechanisms usually introduce random noise into analysis results. An increase in noise magnitude corresponds to heightened privacy protection since it better obfuscates the contribution of each individual. However, the guarantee of privacy comes at the cost of a reduction in accuracy, and the result deviates from the ground truth due to the injection of noise. Many researchers evaluate, estimate, visualize, and optimize the significant accuracy-privacy trade-off in DP [101, 102, 103, 104]. Secondly, the trade-off including performance is usually related to another sort of privatization technique related to cryptography. Although it guarantees both privacy and accuracy by encrypting the data of each individual before sharing it and obtaining the real result by corresponding decryption, there is an enormous amount of computation and communication overhead when performing such operations in FA. The trade-offs among privacy, communication, and computation are adequately investigated by researchers in the cryptography society [105, 106, 107]. Lastly, there are approximating methods applied in FA to increase the performance of the system which sacrifices the accuracy of results. For example, private sampling algorithms (See Section V-C for details) utilize less data in the computation and communication process while introducing variance from sampling into the final result. To conclude, balancing these trade-offs is a key challenge in FA. Researchers in [77] treat the multi-dimensional trade-offs as a major design focus in federated query, and such trade-offs are also taken into great account in many FA solutions. How to properly handle the trade-offs when utilizing the diverse privatization tools in FA that fit the characteristics of data analytics scenarios and host task requirements, becomes a critical challenge in FA design.
IV-C Intensive computation
@@
In FA, cryptography is commonly used to protect the privacy of each client’s data while it is processed and analyzed by multiple parties including the aggregator and other clients. However, cryptographic methods are considerably resource-intensive, both in terms of computation and communication. Operations in cryptography like encryption, decryption, and hashing are computationally intensive. For some advanced cryptographic techniques used in FA like homomorphic encryption and MPC which allow various computations on the encrypted data, the required computation resource is even higher [108]. Regarding communication cost, encrypted data tends to be larger than their plaintext counterparts. Furthermore, most cryptographic protocols require multiple rounds of communication between parties [109], exchanging messages like intermediate results. This attribute of cryptography extremely increases the bandwidth required for tasks in FA. In addition, managing cryptographic keys is also a complex task in federated settings. They are required to be private and secure in their generation, distribution, rotation, and destruction [110]. This process will be more complex and computationally intensive if the number of clients increases [37]. These attributes together contribute to the large resource demands for utilizing cryptography in FA, which is a challenge to preserve privacy efficiently in terms of computation costs. In conclusion, intensive computation is a challenge in many FA solutions, especially those utilizing privatization methodology based on cryptography, according to our taxonomy. Meanwhile, DP-based solutions usually do not introduce a significant computation burden.
IV-D Wireless communication
@@
Wireless communication becomes a major challenge in FA primarily due to its distributed and dynamic nature. The slowest participant in this networked system can significantly hinder overall performance, known as the “straggler effect”. This is often exacerbated by varying data transmission rates and unreliable communication channels, leading to delays and inconsistencies in data aggregation. The paper [111] addresses it in FL by clustering user equipment based on upload times, thereby reducing time divergence among participants. In addition, the over-the-air [112] technique brings efficiency as well as new challenges to FA and FL, such as signal interference and the requirement for precise synchronization. Furthermore, the mobility of devices in wireless networks, such as in vehicular networks and aerial networks, introduces another layer of complexity, requiring adaptive strategies for data transmission and model training that can cope with changing network typologies and varying channel conditions. The integration of technologies like Reconfigurable Intelligent Surfaces shows promise in enhancing communication efficiency and reliability in these scenarios [113], while they also require sophisticated optimization of resources and careful handling of estimated Channel State Information.
IV-E Limited privacy resource
@@
Privacy resource in the context of FA usually refers to the “privacy budget” in the DP mechanisms[114], which intuitively means the amount of allowable privacy loss for each individual within an analysis of the whole data set. Although the information of clients is protected by various privacy-preserving techniques, there is still a risk that repeated analyses such as multiple or iterative queries can expose it. Typically, a portion of the privacy budget is consumed and a specific amount of noise will be added to the results for a query. Managing the privacy budget is important since multiple analyses may be performed on the same group of clients. Every analysis incrementally raises the risk of privacy loss so this cumulative effect ought to be monitored and controlled. Effectively scheduling the privacy budget for the multi-query analysis is a challenge in FA concerning the accuracy of results for each query and total privacy loss for each client. The idea of treating privacy as an allocatable resource is widely accepted, and many prior works present privacy budgeting mechanisms in federated computation systems with a focus on efficiency or fairness [115, 116, 117, 118, 119, 114]. In contrast to the challenge of intensive computation which usually happens in cryptography-based FA solutions, the challenge of limited privacy resource are usually faced by FA solutions with DP-based privatization methodology according to our taxonomy.
IV-F Data heterogeneity
@@
Data heterogeneity, i.e., the phenomenon that data possessed by different clients follows heterogeneous distributions, is common in federated systems. As the tenet of privacy preservation restricts raw data from being transmitted, data reallocation or adjustment is usually unavailable for federated systems, making data heterogeneity unavoidable. The challenge of data heterogeneity has become one of the most popular research focuses in the field of FL [120, 92, 121, 122, 123, 124], and also widely exists in FA systems. Since FA systems tackle various data analytics tasks, data heterogeneity introduces diverse influences for different heterogeneous tasks. In FL applications, data heterogeneity is usually considered harmful to the convergence and accuracy of FL models. In [27], data heterogeneity is studied in a federated clustering problem, where the data points held by one client are biased to one or several clusters. The authors of [27] found that data heterogeneity is beneficial for the FA scheme, where a higher analytics performance can be achieved in the heterogeneous data environment. In [14], an FA scheme is proposed to measure the severity of data heterogeneity directly. The proposed scheme can figure out clients with low/high data heterogeneity, which provides significant information to other federated tasks in handling the data heterogeneity challenges.
@@ V Key Techniques in Federated Analytics
In this section, we investigate the key enabling techniques that are frequently utilized by FA. These key techniques are categorized by their utilization in FA: In Section V-A, the privatization techniques, i.e., those enhance the privacy preservation of FA, are discussed; in Section V-B, the analytics techniques, i.e., that help derive the analytics results, are presented; in Section V-C, the deployment and optimization techniques, which improves the system availability, are presented.
@@ V-A Privatization techniques
The tenet of FA prevents the exposure of raw data to any entity other than the data owner (client), which naturally provides privacy preservation to some extent. However, such a “hiding raw data” idea fails to provide any formal privacy guarantee. Therefore, existing FA research prefers to enhance the privacy preservation of their schemes by providing some formal and rigid privacy guarantees, by leveraging various privatization techniques. In this part, we analyze the privacy guarantees and the corresponding techniques to realize these techniques. Since these privatization techniques can usually be applied in arbitrary computation procedures, the following techniques have gained success in various distributed/centralized computation schemes before FA. However, we would like to emphasize that these techniques are uniquely applied in the novel mechanisms and algorithms in FA, compared to other applications. In the following, we will emphasize the novel usage of these techniques in the field of FA. We would like to emphasize that the privatization techniques of FA are highly related to our taxonomy of privatization methodology. When FA designer decides to apply a particular privatization methodology (like DP or cryptography tools), it becomes necessary to apply the corresponding techniques to enforce the desired privatization methodology.
TABLE V: Comparison and drawbacks of privatization techniques. The multidimensional performance of privatization techniques is marked as good ($+$), medium (++), and bad (+). A privatization technique is considered as good computation/communication efficiency when the execution of such technique introduces small computation/communication overhead.
Technique Privacy Comput. efficiency Commun. efficiency Data utility Scalability Major drawback
     CDP + $+$ $+$ ++ $+$ Assumption of trusted aggregator, limitation in local privacy model
     LDP $+$ $+$ $+$ + $+$ High magnitude noise, loss in data utility
     DDP $+$ ++ ++ ++ ++ Extra overheads from cryptography tools
     $k$-anon. ++ $+$ ++ ++ ++ Loss in data utility, limitation in local privacy
     HE $+$ ++ $+$ $+$ ++ High computation overhead, limited operation support
     MPC $+$ $+$ ++ $+$ + High communication overhead, limited scalability
TABLE VI: Summery of applications of key techniques.
Technique Task Reference
LDP Mean computation [31, 32, 33]
Medians and percentiles [51]
Graph metrics [48, 49, 50]
Frequency-related apps [25, 26, 65]
CDP Mean computation [37, 42, 43]
Frequency-related apps [28, 70]
DDP Frequency-related apps [71, 17, 72]
Database operations [77, 81, 82]
$k$-anonymity Database operations [74, 30]
Assisting federated learning [125]
HE Mean computation [37]
Graph metrics [82]
Database operations [81, 83]
MPC Mean computation [37]
Medians and percentiles [40, 41]
Frequency-related apps [71]
Database operations [75, 78, 79, 80, 81, 83]
Sketching Mean computation [65]
Frequency-related apps [25, 38, 71]
Specialized data structure Mean computation [37]
Frequency-related apps [28, 70, 67]
Optimization theory Database operations [126]
Assisting federated learning [23, 127, 128]
Game theory and incentive design Frequecy-related apps [129]
Database operations [130]
Assisting federated learning [131]
Sampling Frequency-related apps [62, 68]
Database operations [77, 83]
V-A1 Local differential privacy
@@
DP, introduced by [132], is a privatization technique enforced on any data publication mechanism. It has become the most popular criterion for privacy preservation in data analytics, as well as FA. It considers the threat models where the data publication mechanism publishes some outputs based on the raw data, and an adversary tries to infer sensitive information based on the output, such as membership inference attacks [133] and data reconstruction attacks [134]. DP handles such threat models by introducing randomness to the output of the data publication mechanism and restricting the probabilistic distribution of outputs.
@@
LDP, first introduced in [135] [^3] This paper is published even earlier than DP, and it gives a mathematical definition equivalent to LDP., is the most widely used DP variation in FA, as it provides the strong local privacy preservation required by the majority of FA applications. LDP-based FA schemes consider the threat model that the clients upload their insights to an untrusted server, where the untrusted server directly leverages the received insights to perform inferences. LDP is enforced by applying randomness in client uploads so that the probability of any output would not change intensely for clients with arbitrarily different local data. The definition of LDP in an FA setting is given as follows.
@@
The satisfaction of LDP requires perturbation (randomness) on the client uploads. For the binary data, LDP is typically realized by the randomized response, i.e., flipping each bit with some probability. The randomized response-based LDP mechanisms usually work in FA applications where the local computation results are naturally in the binary form [31, 26, 62, 64] or the client data are encoded into a binary form with sketching techniques [25, 66, 65]. For the continuous scaler data, LDP is typically realized by introducing continuous scaler noise on the data. Gaussian mechanism and Laplace mechanism are the most widely used schemes to realize LDP for continuous scaler data. The Gaussian mechanism is applied by adding a random noise following the Gaussian distribution into the uploaded scaler, which can guarantee the corresponding $(\epsilon,\delta)$-LDP [45, 46]. Laplace mechanism, introducing a random noise following the Laplace distribution, can satisfy the stronger $\epsilon$-LDP, which have been utilized in [33, 136, 83].
LDP has been applied in numerous privacy-preserving algorithms other than FA. When applied in FA, many novel utilizations are present. In conventional non-FA applications, LDP noise is usually directly applied to raw data. In FA applications, LDP is in contrary applied on the insight uploaded to the server. In many FA designs, the insights are designed to have a much smaller dimension compared to the raw data, which decreases the magnitude of LDP noise. These examples include the bloom filter in [25], count mean sketch [65], and even one-bit insight [26].
Lessons learned. DP mechanisms provide statistical privacy preservation by adding noise to the client uploads so that the original values cannot be accurately recognized by an adversary. LDP is the DP variation that provides the strongest statistical guarantee. FA considers an adversarial environment without any trusted aggregator, where each client adds significant noise to their uploads to protect their record against any potential aggregator, which fits most of the FA settings. LDP is realized by requiring clients to individually add sufficient noise so that the noise upload cannot reveal sensitive information about the original values. While LDP provides an advantage in strong privacy preservation, its major drawback is the significant magnitude of noise required to satisfy LDP, which may ruin the data utility and limit the design of the insight structure.
V-A2 Central differential privacy
CDP, also known as global differential privacy, is the original form of DP. It considers the scenario that a data analytics mechanism processes a database containing many data samples, and the published analytics results can protect the raw data of any single data sample from being inferred by an adversary. The definition of CDP is as follows.